A COIN IS TOSSED 6 TIMES .FIND THE NO OF GETTING 4 HEADS AND 2 TAILS

Grade:11

1 Answers

Arnab Roy
7 years ago
Use Binomial Theorem nCr (p)^r (q)^n-r probability of getting a head is 1/2=p probability of NOT getting a head is 1/2=q n=6 r=2 6C2 *(1/2)^2 * (1/2)^6-2 = 6C2*(1/2)^6 =15/64